OpenStudy (anonymous):

A landscaper is designing a flower garden in the shape of a trapezoid. She wants the length of the shorter base to be 3 yards greater than the height, and the length of the longer base to be 5 yards greater than the height. For which height will the garden have an area of 225 square yards? Round to the nearest tenth of a yard. >31.3 yards >26.3 yards >121.1 yards >13.1 yards

OpenStudy (anonymous):



OpenStudy (anonymous):

Area of trapezoid = \[\frac{1}{2}*(a+b)*h\]

OpenStudy (anonymous):

Therefore , \[225 = \frac{(x+3+x+5)}{2}*x\]\[225 = \frac{2x+8}{2}*x\]\[225 = \frac{2(x+4)}{2}*x\]\[225 = x(x+4)\]\[225 = x^{2}+4x\]So,\[x^{2}+4x-225 = 0\]Now solve this..